I have a question that states: find the standard form of the function f(x)=x^2-4x+6. Help me please!?!
Answer by stanbon(75887)

find the standard form of the function f(x)=x^2-4x+6.
Assuming you want the form (x-h)^2=4p(y-k)
x^2-4x = y-6
complete the square
x^2-4x+4 = y-6+4
(x-2)^2 = 4(1/4)(y-2)
===================
Cheers,
Stan H.